A box is thrown horizontally from a cliff with an initial velocity of 16m/s, and reaches the ground below in 3.0 s. Find (a) the horizontal distance between the base of the cliff and the location where the box landed and (b) the height of the cliff. (Neglect air resistance.) Use table to list known and unknown variables.

X Y
Xo = 0 Yo =
Δ x = Δ y =
Vox = Voy =
ax =

ay=

t =
Answer #1

Known values are:

vx = 16 m/s

vy = 0 m/s ( box is thrown horizontally)

t = 3 seconds

g = 9.8 m/s2

ax = 0 m/s2

ay = g = 9.8 m/s2

----------------------------------------------------------------------------------------

(a) Use range equation

x = vxt

x = 16 * 3

x = 48 m

This is the horizontal distance between the base of the cliff and the location where the box landed

------------------------------------------------------------------------------------

(b) Use kinematics

h = vyt + 1/2ayt2

h = 0 + 1/2gt2

h = 1/2 * 9.8 * 32

h = 44.1 m

This is the height of the cliff